Bargain Rental Car offers rental cars in an off-airport location near a major tourist destination in California. Management would like to better understand the variable and fixed portions of it car washing costs. The company operates its own car wash facility in which each rental car that is returned is thoroughly cleaned before being released for rental to another customer. Management believes that the variable portion of its car washing costs relates to the number of rental returns. Accordingly, the following data have been compiled: Month Rental Returns Car Wash Costs January 2,400 $ 11,100 February 2,500 $ 13,200 March 2,700 $ 11,900 April 3,000 $ 14,400 May 3,600 $ 16,300 June 5,100 $ 23,500 July 5,500 $ 22,300 August 5,500 $ 22,400 September 4,700 $ 22,900 October 4,000 $ 21,400 November 2,200 $ 10,800 December 2,800 $ 14,600 Exercise 5A-2 Part 2 2. Using least-squares regression, estimate the variable cost per rental return and the monthly fixed cost incurred to wash cars. (Round Fixed cost to the nearest whole dollar amount and the Variable cost per unit to 2 decimal places.)
Answer #1

Answer :

Unit variable cost =

n * total of xy - (Total of x * total of y)

n * total of x^2 -(Total of x)^2

Where

n = 12, Total of xy = 728840000, Total x * total y = 42600 * 187700

Total of x^2 = 167720000 , (Total of x)^2 = 42600 * 42600

Unit variable cost = ((12*728840000) - (42600 * 87700))/((12*167720000 - 1814760000)

=(8746080000 -7996020000) / (2012640000 - 1814760000)

= 750060000 / 197880000

= 3.79

Variable cost per rental return = $3.79

Fixed cost = (Total of y - unit cost * toal of x)/n

=( 187700 - 3.79*42600)/12

=

26246/12

=

$2187

Workings:

Month

Rental returns

Car wash costs

x

y

xy

x^2

January

2300

10200

23460000

5290000

February

2500

12700

31750000

6250000

March

2700

11000

29700000

7290000

April

2900

13100

37990000

8410000

May

3600

15400

55440000

12960000

June

4900

21700

106330000

24010000

July

5400

21400

115560000

29160000

August

5300

20200

107060000

28090000

September

4600

22000

101200000

21160000

October

3800

18700

71060000

14440000

November

2100

9900

20790000

4410000

December

2500

11400

28500000

6250000

Total

42600

187700

728840000

167720000